### v1

@@ 7.1 A simulation comparison of the set estimators of overall ranking
To run a simulation for a setup similar to the baseball example we generate 18 values of a covariate independently from a uniform (0,1) distribution. These covariates are kept fixed throughout the simulation. Using them, we generate $\theta_{i}$’s independently from $N(\beta_{0}+\beta_{1}x_{i},A)$. Using the generated $\theta_{i}$, we generate $Y_{i}$ from $N(\theta_{i},D_{i})$. We use the $D_{i}$ values from the baseball application. We use five possible values of $A=0.001,0.005,0.01,0.1$ and $1$ (for the baseball data estimate of $A$ is about $0.002$). We use $\beta_{1}=0$ (no covariate) and $\beta_{1}=0.4$. We use $\beta_{0}=0.2$, although $\beta_{0}$ does not influence the ranks and the results.
We report our simulation results in Table 3. First row of each block of the table, with a W in parentheses, reports average posterior expected deviations for weighted credible distribution, and the second row with UW in parentheses, shows results for its unweighted version. In the left half of these tables we presented results with a covariate and in the right half with no covariate. For each parameter setting, we provided the average of the mean absolute deviations of the overall ranks from the true ranks for KWW method, and average of the posterior expected absolute deviations for Bayesian methods: HB and UB. We also presented average volumes and average of lengths of the sides of various 90% confidence or credible sets of the true mean vector $\theta$. When we compare in terms of size measures, the $m$th root of the volume (to avoid dealing with very large or very small numbers, we consider the $m$th root) and the average length of the confidence or credible sets of $\theta$, the HB elliptical credible set performed the best. In terms of the average posterior expected deviations of ranks from their true values, the performance of the HB solution is mixed and the Klein et al. (2020) method has the weakest performance. Smaller values of $A$ (0.001, 0.005, 0.01), relative to the $D_{i}$’s, introduce a lot of structure to the $\theta_{i}$’s. In these cases in the absence of any covariate, $\theta_{i}$’s tend to be a lot similar, where the UB method performs better based on deviation criterion. However, when the $\theta_{i}$’s vary according to a covariate, the HB method performs substantially better for smaller values of $A$. For larger values of $A$, $0.1$ and $1$, even with the covariate, $\theta_{i}$’s do not display much of any structure (relative to the $D_{i}$’s, $A=0.1$ is between 19 and 35 times the $D_{i}$’s, and $A=1$ is between 190 and 350 times larger). The case of $A=1$ relative to the $D_{i}$’s is similar to our next example where we develop ranking of the US states based on average commuting times. For the last two cases of $A$, the HB posterior distribution of $\theta$ will have negligible shrinkage and it is approximately equal to unstructured Bayesian posterior distribution, resulting in similar inference for credible distribution of overall ranking; the UB method is marginally better than the HB method in terms of the deviation measure. Finally, in terms of this measure, the overall ranking for the elliptical method does better than the Cartesian method when the double stochastic matrix is created by unweighted method. However, when the credible distribution is created by weighted method, the credible distributions from elliptical and the Cartesian sets perform very similarly.
@@
8 Ranking of the US states based on commuting times
In a pioneering article, Klein et al. (2020) applied their frequentist approach to rank fifty states of the U.S. and DC by mean commuting times of workers sixteen or older and not working from home. They used survey data collected from the ACS. In Figure 2 below, we recreated Figure 1 of Klein et al. (2020) that depicted the frequentist solution of the confidence region for ranking from equation (8). Possible ranks for state $i$ in this set are shown by a pink line segment stretching from $|\Lambda_{Li}|+1$ to $|\Lambda_{Li}|+|\Lambda_{Oi}|+1$. From the figure, for example, the possible ranks from this solution for the state ID are $4-9$. On the other hand, the states which can hold the rank 9 are the states WY, AK, MT, IA, KS and ID (the pink line segments for these states intersect the horizontal line for ranks = 9). The rank based on direct estimate $y_{i}$ is shown by a black dot at its rank $r$. If two states tie for ranks $r,r+1$, their tied ranks are shown by a line segment joining these states at rank $r+1$. For example, SD and ND direct estimates tie for ranks 1 and 2. We showed this by the black line joining the two states.
Figure 2: An alternative visualization of the 90% joint confidence region for travel time ranking given by Figure 1 of Klein et al. (2020) with an overlapping credible distribution from unstructured Bayes method
On Figure 2 we overlaid the credible distribution based on a Cartesian credible set from the unstructured Bayes model. Both in terms of the model and the method, this table is most directly comparable with Figure 1 of Klein et al. (2020). Probabilities from credible distribution are shown by yellow circles, bigger circles for larger probabilities. To clearly interpret the probabilities depicted in the figure, let us focus specifically on its $4$th row and $4$th column. From the $4$th column of this figure we find that the state of NE can have ranks 3-6 with respective probabilities about $0.2,0.2,0.5,0.1$ (actual values are $0.21,0.20,0.55,0.04$). We can contrast this with the red band on the 4th column in the figure which shows the rank set for NE suggested by the Klein et al. (2020) solution. It shows that in addition to the four ranks 3-6, NE can also rank 1, 2, 7 and 8. On the other hand, the fourth row of the figure corresponding to rank 4 shows that the states that can plausibly hold this rank are AK, MT, NE and WY with respective probabilities about $0.05,0.2,0.5,0.2$ (actual values are $0.04,0.23,0.55$ and $0.18$). Again, in contrast, the 4th row in Figure 1 of Klein et al. (2020) shows that states SD, ND, WY, NE, MT, AK, IA, KS and ID are all contenders of rank 4. Compared to the Bayesian solution, the frequentist solution includes five additional states. From all the posterior samples of the vector $\theta$ (not only those that are included in the credible set), in Figure 3 we created the scatter plot of NE vs. IA, using their components from $\theta$. The entire scatter plot is practically below the line $y=x$, implying that it is improbable for IA and NE to hold the same rank. As another example, if we consider ID, we find from our table that it can hold only rank 9, and the rank 9 can only be held by Idaho. On the other hand, we observed it earlier that for the frequentist solution possible ranks for ID are 4-9, and the states that can rank 9th are WY, MT, AK, IA, KS and ID. Similarly, while SD and ND, tied by direct estimates, and both have the same frequentist rank set $1-6$, according to credible distribution ranks, only possible ranks for these states by the credible distribution are $1$ and $2$, and SD has $60\%$ probability to hold rank $1$ (note that SD direct estimate has a smaller standard error). Clearly, all these results demonstrate that our Bayesian results are “less dispersed”, “more illuminating” and “more conclusive”.
Figure 3: Scatter plot of the posterior samples of commuting times for IA and NE
We also compute the credible distribution (i.e., the double stochastic matrix), not shown here, based on the elliptical region from the HB model. From the $4$th row of this matrix we find that the states AK, MT, NE and WY are the contenders of rank 4, and they may hold this rank with probabilities $0.11,0.21,0.50$ and $0.18$, respectively. The fourth column of that matrix conveys that NE can hold the ranks 3, 4, 5 and 6 with probabilities $0.32,0.50,0.15$ and $0.04$, respectively. Close agreements between these probabilities and those for the other cases show that credible regions of $\theta$ having different shapes and derived from different models provide reasonably similar solutions for overall ranking. We have no covariate for this example and the HB estimate of the model error variance $A$ is $13.2$, which is about $56$ times the largest sampling variance $D_{i}$ and $3572$ times the smallest one. With so high ratios of the model variance to sampling variances, the HB model is practically identical to the UB model. Similar credible distributions from the UB and the HB models in this application are confirmed by the simulation study we presented in Table 3. From the two lower-right-bottom blocks of this table, we find the average posterior mean absolute deviations reported for the UB and the HB columns show no practical differences among the four Bayesian solutions (2 models times 2 shapes). In the last block we set $A=1$ which is nearly $188$ times the largest sampling variance and $342$ times the smallest one.
We also compare all five confidence and credible regions of $\theta$ in terms of their size measures. From Table 4 we find that all four Bayesian credible sets have smaller volumes than the KWW confidence set. While volumes of Cartesian credible sets are between $31\%$ and $38\%$ smaller than the volume of the KWW set, the elliptical shape HPD credible sets are enormously superior (by an order of $10^{-7}$) to the other set estimators with hyperrectangular shapes (this is due to cutting off low probability “corners” of the hyperrectangular sets to create an HPD set). Even in terms of average length, elliptical-shaped sets are better than the hyperrectangular-shaped sets. Finally, with no surprise to us, we observe that in terms of these size measures credible sets from the HB model are better than the credible sets from the unstructured Bayes model.
Table 4: Commuting time: Comparison of size of set estimators of $\theta$, frequentist and Bayes.
Frequentist Unstructured Bayes Hierarchical Bayes
Measure KWW Cartesian Elliptical Cartesian Elliptical
Volume 4.52 3.54 $1.89\times 10^{-7}$ 3.42 $1.46\times 10^{-7}$
Ave length 1.16 1.16 0.83 1.15 0.83
9 Ranking of the US states based on median incomes
Inference on median incomes of the US states for the income year 1989 has served as a benchmark example to evaluate effectiveness of various small area estimation methods. See for example, Ghosh et al. (1996) and Chung and Datta (2022). For this application, these researchers suggested various innovative small area estimation methods to estimate the true four-person family median incomes ($\theta_{i}$) for the fifty US states and Washington, DC. They used survey data from the Current Population Survey (CPS) that provides annual estimates of median incomes for the states for various family sizes. These estimates serve as the direct estimates for the popular Fay-Herriot model. The application enables borrowing strength from other covariates by integrating administrative data and the past census data. For this application Fay (1987) suggested in addition to the intercept using two covariates in the Fay-Herriot model in Section 2.2. The first covariate $x_{i1}$ is the $i$th state median income for 1979 from the preceding Census in 1980. The second covariate, $x_{i2}$, is an adjusted version of $x_{i1}$. The adjusted Census median income $x_{i2}$ for the $i$th state is given by $x_{i2}=(\text{PCI}_{i,1989}/\text{PCI}_{i,1979})x_{i1}$, $i=1,\ldots,m$, where $\text{PCI}_{i,1979}$ and $\text{PCI}_{i,1989}$ are the 1979 and 1989 per capita incomes of the $i$th state provided by the Bureau of Economic Analysis of the U.S. Department of Commerce. Small area estimates using these covariates have been effectively compared with the corresponding incomes obtained from the 1990 Census that collected data for the 1989 income year. The census incomes for the states were medians based on a large number of households from each state. Due to large sample sizes the state census median incomes are treated as “gold standards” and are believed to be effective “surrogates” for the true 1989 state median incomes, $\theta_{i}$, which we are trying to rank.
Figure 4: A visualization of the 90% joint confidence region for median income ranking by Klein et al. (2020) with an overlapping credible distribution from the hierarchical Bayes method
In Figure 4, we recreated for our application Figure 1 of Klein et al. (2020) that depicted the frequentist solution of the confidence region for ranking from equation (8). On this figure we overlaid the credible distribution based on the weighted HB model. The KWW rank sets for all the states are depicted by appropriate yellow line segments that span a subset of 1 to 51. Probabilities from credible distribution are shown by red circles, bigger circles for larger probabilities. We use solid black circles on this figure to depict the “true” states ranks based on their median incomes from the 1990 census by. For our application, the frequentist KWW solution assigns ranks 1 to 46 to the state MS; it is hardly informative. The HB credible distribution assigns probabilities $0.25,0.11,0.14,0.14,0.06,$ and $0.20$ to the ranks 1 to 6. Rest of the ranks have negligible probabilities, adding up to $0.1$. To the state CT the KWW solution assigns ranks 15 to 51, which is again very wide. The HB Cartesian credible distribution assigns probabilities $0.03,0.13,0.33,$ and $0.51$ to the ranks 48 to 51. Clearly, the Bayesian solution is very compact. For NJ, the KWW ranks span from 22 to 51. The HB solution assigns probabilities $0.04,0.53,$ and $0.43$ to the ranks 49-51. The Bayesian solution is vastly superior to the frequentist solution.
Table 5: Median incomes: Comparison of set estimators of $\theta$, frequentist and Bayes, by size measures, and comparison of overall ranking methods by total average absolute deviations from gold standard ranks. For each of the two Bayesian models, we have two methods based on elliptical (subscript E) and Cartesian (subscript C) sets. All volume measure have been scaled by a known constant $g$ and all length measures have been scaled by a known constant $h$.
Measures $\epsilon_{i,\small{HB,C}}$ $\epsilon_{i,HB,E}$ $\epsilon_{i,UB,C}$ $\epsilon_{i,UB,E}$ $\epsilon_{i,KWW}$
Total abs. dev. 179.52 179.52 388.11 388.11 798.66
Volume/$g$ $4.619\times 10^{19}$ $2.975\times 10^{12}$ $8.321\times 10^{26}$ $4.043\times 10^{19}$ $1.080\times 10^{27}$

### v2

@@ 10.1 A simulation comparing set estimators of overall ranking
To run a simulation for a setup similar to the baseball example we generate 18 $x_{i}$’s for a covariate independently from a uniform (0,1) distribution. These covariates are kept fixed throughout the simulation. Using them, we generate $\theta_{i}$’s independently from $N(\beta_{0}+\beta_{1}x_{i},A)$. Using the generated $\theta_{i}$, we generate $Y_{i}$ from $N(\theta_{i},D_{i})$. We use the $D_{i}$ values from the baseball application. We use five possible values of $A=0.001,0.005,0.01,0.1$ and $1$ (for the baseball data estimate of $A$ is about $0.002$). We use $\beta_{1}=0$ (no covariate) and $\beta_{1}=0.4$. We use $\beta_{0}=0.2$, although $\beta_{0}$ does not influence the ranks and the results.
We report our simulation results in Table 4. First row of each block of the table, with a W in parentheses, reports average posterior expected deviations for weighted credible distribution, and the second row with UW in parentheses, shows results for its unweighted version. In the left half of these tables we presented results with a covariate and in the right half with no covariate. We also provided the average of the mean absolute deviations of the overall ranks from the true ranks for KWW method, and average of the posterior expected absolute deviations for Bayesian methods: HB and UB. In the last two rows of each block, we presented average volumes and average of lengths of the sides of various 90% confidence or credible sets of the true mean vector $\theta$.
@@
When compared in terms of size measures, the $m$th root of the volume (to avoid dealing with very large or very small numbers) and the average length of the confidence or credible sets of $\theta$, the HB elliptical credible set performed the best. In terms of the average posterior expected deviations of ranks from their true values, the performance of the HB solution is mixed and the Klein et al. (2020) method has the weakest performance. Smaller values of $A$ (0.001, 0.005, 0.01), relative to the $D_{i}$’s, introduce a lot of structure to the $\theta_{i}$’s. In these cases in the absence of any covariate, $\theta_{i}$’s tend to be a lot similar, where the UB method performs better based on deviation criterion. However, when the $\theta_{i}$’s vary according to a covariate, the HB method performs substantially better for smaller values of $A$. For larger values of $A$, $0.1$ and $1$, even with the covariate, $\theta_{i}$’s do not display much of any structure (relative to the $D_{i}$’s, $A=0.1$ is between 19 and 35 times the $D_{i}$’s, and $A=1$ is between 190 and 350 times larger). The case of $A=1$ relative to the $D_{i}$’s is similar to our example where we develop ranking of the US states based on average commuting times. For the last two cases of $A$, the HB posterior distribution of $\theta$ will have negligible shrinkage and it is approximately equal to UB posterior distribution, resulting in similar CDs for overall ranking; the UB method is marginally better than the HB method in terms of the deviation measure. Finally, in terms of this measure, the overall ranking for the elliptical method does better than the Cartesian method when the double stochastic matrix is created by unweighted method. However, when the CDs are created by weighted method, the CDs from elliptical and the Cartesian sets perform very similarly.
11 Summary and Conclusions
Inference for population means (or proportions) of subpopulations or subjects is an important statistical problem. Estimation of (i) small area characteristics in survey sampling, (ii) breeding values of various herds in animal science and livestock research, (iii) true performance of various players in sports, and (iv) treatment effectiveness for certain conditions treated by several health facilities in healthcare, for example, are some of the important problems in multi-population studies. The major focus in statistical inference for multiple populations has been estimation (point and interval) of means or proportions or similar measures from several subpopulations or individuals. In small area estimation, the companion problem of simultaneously estimating the ranks of the small areas is also of considerable importance but it has not gotten enough attention. Only Shen and Louis (1998) simultaneously considered estimation of small area means and ranks. Using the HB approach these authors considered point estimation of ranks but did not account for error in estimation of the rank vector for all small areas.
To account for estimation error for the rank vector of many subpopulations, in a pioneering article Klein et al. (2020) considered set estimation of overall ranking. Klein et al. (2020) created joint confidence set of the rank vector by employing a rectangular confidence set of the mean vector $\theta$. Their solution in equation (8) attributes $|\Lambda_{Li}|+1,\cdots,|\Lambda_{Li}|+|\Lambda_{Oi}|+1|$ as the possible values for the true rank ${\check{r}}_{i},i=1,\cdots,m.$ This solution, frequentist by construction, may be interpreted as a confidence distribution for ${\check{r}}_{i}$, uniformly distributed over the integers between the lower and the upper bounds presented above. As an alternative to this method, which cannot use auxiliary information or model the unknown true mean $\theta$, we introduced a Bayesian model for the mean vector. The model allows us to utilize useful auxiliary information as well as to develop various credible sets for $\theta$ including the HPD sets that are exact or nearly exact. We created various credible distributions for the true rank ${\check{r}}_{i}$, which turn out to be more informative than the confidence set by KWW. In our Bayesian proposal we can also allow more probable $\theta$’s (measured by the posterior pdf) in the credible set to have them bigger weights to the credible distributions.
Application of proposed CDs is computationally straightforward. We applied our solution to three applications, two to rank small areas (discussed extensively) and one to rank professional baseball players (reported briefly accompanied by extensive discussion in the Supplement). All three applications and a simulation study mimicking two of the applications demonstrated superior performance of the proposed CDs over the existing frequentist method.
For the baseball example there are no reliable covariates but there are surrogate values of the true means from the players’ performance from the remainder of the season. Evaluation of frequentist confidence set and the credible distributions based on the average or expected absolute deviations of estimated ranks from the ranks of the surrogate values show better performance for the Bayesian solutions. In the absence of covariates, the UB credible distributions perform better than the HB credible distributions. For the commuting time example we neither have any covariates nor any representative values of the unknown true means. Even then in this example the credible distributions are “more conclusive”, “less dispersed” and “more illuminating” than the frequentist solution by Klein et al. (2020). For the median income estimation problem we have useful covariates as well as representative values (from the 1990 census) of the true means. Credible distributions from the HB model with the covariates performed the best when measured in terms of average absolute deviations. Both credible distributions, from the UB model and the HB model, have better performance than the frequentist method. The HB credible distributions based on covariates are uniformly better than the other two solutions. However, the HB credible distributions without any covariates do not enjoy a uniform superiority over the other two solutions. These outcomes are confirmed by our simulation study for the baseball setup. Availability of good covariates shows the superiority of HB credible distributions. Lack of covariates degrades the performance of HB credible distributions, which are sometimes inferior to the UB credible distributions. In terms of size measures of the confidence or credible sets, the Bayesian solutions are vastly superior, and between the two Bayesian credible sets, the HB-based sets are shorter or more “compact” than the UB-based sets.
Supplementary Materials
S.1 Gibbs sampling: The set of full conditional pdfs for the Gibbs sampling.
S.2 Size measures of elliptical sets: Discussion of two size measures of elliptical sets.
S.3 More results on commuting and income data: A double scatter plot, Figure S.1, from the HB posterior sample of commuting times discussed in Section 8, an extraction of the double stochastic matrix, Table S.1, from the same posterior sample, and Table S.2, a table of average or expected absolute deviations of ranking from the surrogate median incomes discussed in Section 9.
S.4 Ranking of baseball player: A full discussion of the application ranking of American major league professional baseball batters, discussed briefly in Section 10. This example was considered by Efron and Morris (1975) demonstrating the usefulness of EB inference.
Disclosure Statement: No competing interests to declare from the authors.
Funding
Supplementary Materials
S.1 The set of full conditional pdfs for the Gibbs sampling
The joint posterior pdf of $\theta,\beta,A$ is
$\pi(\theta,\beta,A|y)$ $\propto$ $\exp[-\frac{1}{2}(y-\theta)^{T}D^{-1}(y-\theta)]\times A^{-\frac{m}{2}}\exp[-\frac{1}{2A}(\theta-X\beta)^{T}(\theta-X\beta)]$
$\times$ $(\bar{D}+A)^{-\frac{1}{2}}.$
From the above the full conditionals are
  • (A)
    $\theta|\beta,A,y\sim N(\delta,\Phi),$
  • (B)
    $\beta|\theta,A,y\sim N((X^{T}X)^{-1}X^{T}\theta,A(X^{T}X)^{-1})$,
  • (C)
    $\pi(A|\theta,\beta,y)\propto(\bar{D}+A)^{-\frac{1}{2}}A^{-\frac{m}{2}}\exp[-\frac{1}{2A}(\theta-X\beta)^{T}(\theta-X\beta)]$,
where
$\Phi=(D^{-1}+A^{-1}I_{m})^{-1},~{}\delta=\Phi(D^{-1}y+A^{-1}X\beta),$
To sample from the conditional pdf of $A$ we use the rejection sampling. We use the inverse gamma distribution with shape parameter $(m-1)/2$ and rate parameter $(\theta-X\beta)^{T}(\theta-X\beta)/2$ as the proposal distribution. Alternatively, one can also use the Metropolis-Hastings algorithm within Gibbs with the same proposal as the transition density for $A$.
S.2 Size measures of elliptical sets
As an alternative to an orthotope (or hyperrectangular) credible set, where each component of $\theta$ varies freely, we have considered HPD sets which are approximately elliptical for our Bayesian models. Two size measures that appear to be appropriate for set estimators are their Lebesgue volumes and average lengths of the “sides” determining the set. These two measures are easily found for an orthotope. We compute representative values of these size measures of the elliptical sets.
To determine reasonable size measures for elliptical regions, we proceed with the set
$E=\{x:x^{T}\Phi x\leq c\},$ (S.1)
where $\Phi$ is an $m\times m$, p.d. matrix, and $c>0$, determine the region. Since the center of the ellipse has no role in the size measures, we take it zero. Let $V_{E}(\Phi,m,c)$ denote the Lebesgue volume of the set $E$. It can be easily calculated that
$V_{E}(\Phi,m,c)=\frac{\pi^{m/2}}{\Gamma((m+2)/2)}\times\frac{c^{m/2}}{\mbox{det}(\Phi)^{1/2}},$ (S.2)
where $\Gamma(\cdot)$ is the gamma function and $\mbox{det}(\Phi)$ is the determinant of $\Phi$.
A standard length measure of an elliptical region is not available. We propose below a reasonable expression for it. For an $m$-dimensional rectangular region, its volume is the product of the volume of the $(m-1)$-dimensional of the cross-section of the set defined by the intersecting hyperplane, say, $\theta_{1}=d$, and the length $(U_{1}-L_{1})$ of original rectangular set along the $\theta_{1}$ direction (where $L_{1}<d<U_{1}$). Thus, the length of the set along the $x_{1}$ direction is the ratio of the volume of the original orthotope to the volume of the lower-dimensional cross-section orthotope.
We will use the above discussion for the orthotope region as an analogy to come up with a suitable “length” for the ellipse along the $x_{1}$ direction. For the ellipse, we take for any $x_{1}$ and $x_{(2)}=(x_{2},\cdots,x_{m})^{T}$ with $(x_{1},x_{(2)}^{T})^{T}$ inside the set, we define a central cross-section by the hyperplane $x_{1}=-\Phi_{12}x_{(2)}/\Phi_{11}$, where we partition the matrix $\Phi$ as $((\Phi_{ij}))$, $i,j=1,2$, in conformity with the partitioning of $x$ above. We also define the p.d. matrix $\Phi_{22.1}=\Phi_{22}-\Phi_{21}\Phi_{11}^{-1}\Phi_{12}$. The corresponding cross-section is $E_{C}=\{x_{(2)}^{T}\Phi_{22.1}x_{(2)}\leq c\}.$ From equation (S.2), it similarly follows that
$V_{E_{C}}(\Phi_{22.1},m-1,c)=\frac{\pi^{(m-1)/2}}{\Gamma((m+1)/2)}\times\frac{c^{(m-1)/2}}{\mbox{det}(\Phi_{22.1})^{1/2}}.$ (S.3)
A reasonably representative length for the $x_{1}$ side of the ellipse is the ratio of the volume in (S.2) to the volume in (S.3). Simplification leads to an expression. From this, after simplification, in general, we have along the $x_{i}$ direction the representative length, $L_{R,i}$ (say), is
$L_{R,i}=B\left(\frac{1}{2},\frac{m+1}{2}\right)\times\left\{\frac{\sqrt{c}}{\sqrt{\Phi_{ii}}}\right\},~{}i=1,\cdots,m,$ (S.4)
where $B(a,b)$ is the regular beta function.
The above proposal of the length of a side of an elliptical region is only heuristic. Indeed it is a reasonable requirement that these lengths should be “calibrated” by a suitable scale so that for the modified orthotope based on the lengths of the “scaled sides” will match the volume of the original ellipse. Based on this consideration, the modified lengths $L_{M,i}$ are given by
$L_{M,i}=\{V_{E}(\Phi,m,c)/\prod_{i=1}^{m}L_{R,i}\}^{1/m}L_{R,i}.$ (S.5)
$L_{E}\stackrel{{\scriptstyle def}}{{=}}\frac{1}{m}\sum_{i=1}^{m}L_{M,i}$, a “linear” measure of the elliptical set, may be compared to $L_{C}$.
S.3 More Analysis for Commuting and Income Data